### v1

# Materialist: Physically Based Editing Using Single-Image Inverse Rendering

Abstract
To perform image editing based on single-view, inverse physically based rendering, we present a method combining a learning-based approach with progressive differentiable rendering. Given an image, our method leverages neural networks to predict initial material properties. Progressive differentiable rendering is then used to optimize the environment map and refine the material properties with the goal of closely matching the rendered result to the input image. We require only a single image while other inverse rendering methods based on the rendering equation require multiple views. In comparison to single-view methods that rely on neural renderers, our approach achieves more realistic light material interactions, accurate shadows, and global illumination. Furthermore, with optimized material properties and illumination, our method enables a variety of tasks, including physically based material editing, object insertion, and relighting. We also propose a method for material transparency editing that operates effectively without requiring full scene geometry. Compared with methods based on Stable Diffusion, our approach offers stronger interpretability and more realistic light refraction based on empirical results. Code is avaiable at github.com/lez-s/Materialist
Figure 1: Given an image (in the middle), our inverse rendering approach enables physically based image editing, including transparency, albedo, roughness, metallic, transparent object insertion and relighting. The image is sourced from the IIW dataset [4].
1 Introduction
High-quality image editing often requires professional skills. Reducing the complexity and increasing the accuracy of image editing has long been a focus in computer vision and computer graphics [36, 40, 45, 3]. With the success of Stable Diffusion (SD) [48] in image generation, researchers have explored using SD for image editing [19, 38, 7, 77]. However, SD often struggles with precise material property editing. Alchemist [53] attempts to address this by training SD on a large synthetic dataset, but their method lacks interpretability, a physical foundation, and struggles with accurate light refraction for transparent objects.
Inverse rendering offers a promising solution to these challenges. Single-view inverse rendering [28, 29, 79, 52] decomposes images into albedo, roughness, and metallic properties via neural networks (MaterialNet), enhancing interpretability. However, relying on a neural renderer as a substitute for a physically based renderer still shares limitations with SD, such as difficulties in accurately handling of shadows and refraction. Multi-view inverse rendering [67, 1, 69, 71, 30, 2] leverages multi-view constraints and the rendering equation to adhere better to physical principles and achieve superior results. The multi-view input requirement, however, restricts its applicability. Applying differentiable rendering to a single image leads to an ill-posed problem with infinitely many possible solutions.
To address these issues, we adopt a more interpretable inverse rendering approach that combines the strengths of single- and multi-view methods (MaterialNet and the rendering equation). Additionally, we introduce material transparency editing, a capability not previously explored in single-image inverse rendering.
Our main contributions are threefold:
  • Hybrid Approach: We integrate learning-based methods with differentiable rendering for a hybrid approach applicable to various inverse rendering tasks. To address the infinite optimization possibilities in single-view differentiable rendering, we use neural network-predicted material properties as pseudo ground truth to guide the optimization process.
  • Physically Based Material Editing: Instead of a neural renderer, our method leverages Mitsuba [20] to achieve physically based material editing with precise global illumination, shadows, and realistic object-environment interactions, thus outperforming existing methods in material editing, relighting, and transparent object insertion tasks.
  • Single-View Transparency Editing: We propose a single-view, physically based transparency editing method with index of refraction (IOR) and specular transmission, which operates under the single-view constraint without requiring complete geometry or full scene reconstruction.
In addition, we demonstrate that our material prediction model outperforms the latest model based on Stable Diffusion [24] with much faster inference. We also achieve better results in single-view differentiable Monte Carlo ray tracing for environment maps and introduce a mesh reconstruction method well-suited for single-image inverse rendering.
2 Related work
Figure 2: Our inverse rendering pipeline. Given an image, we use MatNet to predict material properties, followed by envmap optimization to estimate the lighting. We then perform material properties optimization, using the envmap that yields the smallest $\mathcal{L}_{\text{re}}$ during envmap optimization as the light source. The losses $\mathcal{L}_{\text{re}}$ and $\mathcal{L}_{\text{cons}}$ guide this process, allowing $\mathcal{L}_{\text{re}}$ to be minimized while keeping the results close to the MatNet predictions. See Sec. 3.2 for further details.
Multi-view inverse rendering.
Early work like IPT [1] was limited to simple scenes. They used multi-view constraints and assumed known scene geometry when using the rendering equation for inverse rendering. With NeRF [37], physically based inverse rendering with multi-view images became more feasible [69, 71, 30, 2, 56]. In addition, use of differentiable Monte Carlo rendering led to better shape and material estimation [33, 18, 76, 67, 73, 64], but these techniques still rely on multi-view inputs. More recent methods employ neural networks as renderers for multi-view inputs [12], bypassing computational optimization but currently lacking the performance of equation-based methods in material editing and object insertion. Like most multi-view techniques, our approach uses the rendering equation but requires only a single-view image as input.
Single-view inverse rendering.
Unlike multi-view inverse rendering, which relies on the rendering equation, single-view inverse rendering often uses neural networks trained on large datasets to replace physically based rendering [26, 27, 29, 49, 52, 79, 34]. A neural network usually called MaterialNet, predicts material properties (albedo, roughness, metallic) and per-pixel lighting from a single image, which are then used for image synthesis. Although neural renderers perform well in specific tasks like human face relighting [70, 41], they struggle with generalist tasks where physical simulation is more accurate such as material editing and transparent object insertion. Unless trained on a well represented dataset, the neural renderers fall apart.
In differentiable rendering, early studies addressed rasterization non-differentiability [11, 10], while recent work has explored inverse Monte Carlo ray tracing [21, 75, 14], similar to our physics-based optimization. However, these methods require significant priors and are limited to single objects, making them less suitable for complex scenes. Our approach also uses MatNet for single-view material property inference but incorporates a physics-based renderer.
Material Editing.
Material editing generally follows two categories. The first involves inferring material properties through inverse rendering, then modifying them to achieve edits [25, 78, 1]. These methods provide strong interpretability, and using a physics-based renderer makes results highly reliable, though realism may sometimes be limited. The second set is neural network-based. Following the success of Stable Diffusion [48], many image editing methods have emerged [38, 63, 7, 55]. Alchemist [53] is a representative example, trained on synthetic data to allow editing of albedo, roughness, metallic properties, and transparency. These methods benefit from the strengths of Stable Diffusion and achieve highly realistic results but have limited interpretability due to their neural network-based nature. Our approach falls into the first category with a solid interpretability, meanwhile, it enables realistic results.
Light Estimation.
Light estimation is a distinct research field [62, 60, 42, 57, 16, 72, 58, 13]. Some methods represent lighting implicitly [69, 78, 24], limiting generalizability, while others require multi-view inputs and scene mesh data [60, 72, 42]. DPI [35] is related to our approach. This method combines differential rendering with Stable Diffusion for high-quality envmap generation but relies on multi-view NeRF methods for mesh reconstruction and lacks material BRDF data, reducing accuracy. In contrast, our approach uses MatNet’s output for more accurate envmap optimization.
Shadows and Ray Tracing.
Shadows are crucial for photorealistic rendering. Previous work by Li et al. [29] emphasized shadow accuracy, using an OptiX-based ray tracer [43] to compute shadows for neural network input. However, shadows are often neglected in recent research [78, 69, 30]; methods based on screen-space or image-based ray tracing lack geometric occlusion and thus fail to generate accurate shadows [69, 24, 78]. Many neural renderers struggle with shadow generation [12, 79], as accurate shadows require physically based light transport, it is worth noting that this limitation has been utilized to identify AI-generated images [50]. Our method includes full path tracing for physically accurate shadows.
3 Method
Our material estimation and editing framework consists of four main steps. For relighting, only the geometry reconstruction and the material prediction network (Sec. 3.1) are needed. Object insertion and material editing on the other hand require all four steps, including mesh reconstruction. We assume a pinhole camera model with a 35-degree field of view, $w\times{h}=512\times 512$ image resolution, and $c=3$ color channels.
3.1 Material Prediction Network
Due to the significant success of Stable Diffusion [48] in image generation, recent work used it for the task of material property estimation, achieving impressive results [24]. However, this approach comes with high training costs and requires computationally intensive multi-step sampling during prediction. In comparison, we found that the model architecture for depth estimation [68] aligns more closely with the requirements of material prediction. We therefore adopt the DPT architecture [46], which has been highly successful in depth estimation, for our material prediction task. Different from [79], we utilize a pretrained DINOv2 [39] encoder for feature extraction. Like previous research in depth estimation [6, 68], we use two DPT decoders [46] for the regression of depth and material properties, respectively. Given the advancements in depth estimation, we initialize our training with weights from the pretrained depth model [68] to expedite the training process. For the DPT depth decoder, we limit training to the last four layers of the refinenet, while the weights of the DINOv2 encoder remain frozen throughout the training. The following loss function for $\mathcal{L}_{\text{MatNet}}$ is used to optimize the model,
$\mathcal{L}_{\text{MatNet}}=\sum_{i\in\{\mathbf{A},\mathbf{R},\mathbf{M},\mathbf{N},\mathbf{D}\}}\mathcal{L}_{i}\,,$ (1)
where the albedo loss $\mathcal{L}_{\mathbf{A}}$ combines the LPIPS perceptual loss and $L_{1}$ loss, while both the roughness loss $\mathcal{L}_{\mathbf{R}}$ and metallic loss $\mathcal{L}_{\mathbf{M}}$ are measured using the $L_{1}$ loss between the predictions and ground truth values. Finally, the $\mathcal{L}_{\mathbf{D}}$ is calculated using SiLogLoss [5, 15] to introduce logarithmic error and keep its scale constant to improve the robustness of depth prediction. This loss is defined by
$\mathcal{L}_{\mathbf{D}}=\sqrt{\begin{aligned} &\mathbb{E}_{i\in\text{mask}}\left(\left(\log(t_{i})-\log(p_{i})\right)^{2}\right)\\ &-\lambda\left(\mathbb{E}_{i\in\text{mask}}\left(\log(t_{i})-\log(p_{i})\right)\right)^{2}\,,\end{aligned}}$ (2)
where $t_{i}$ is the ground truth and $p_{i}$ is the predicted depth value for the valid pixel $i$, and we use $\lambda=0.5$ as the hyperparameter that balances the variance and mean terms. To prevent outliers in the depth map from causing the model to collapse during training, we set the mask to a depth value in $[0,20]$. Finally, $\mathcal{L}_{\mathbf{N}}$ is calculated using cosine similarity.
The prediction process of the neural network can be expressed as
$\mathbf{A}_{p},\mathbf{R}_{p},\mathbf{M}_{p},\mathbf{N}_{p},\mathbf{D}_{p}=\text{MatNet}(I)\,,$ (3)
where $\mathbf{A}_{p},\mathbf{R}_{p},\mathbf{M}_{p},\mathbf{N}_{p},\mathbf{D}_{p}$ represent the albedo, roughness, metallic, normal, and depth predicted by MatNet and $I$ is the input image.
Mesh Reconstruction.
Using $\mathbf{D}_{p}$, 2D pixels are projected into 3D to form a point cloud, and neighboring pixels are connected to create a triangle mesh. This method produces a single connected component, which is unsuitable for multi-object scenes.
Figure 3: Our single-view mesh reconstruction method used on a test scene. Our method eliminates the white border at the edges. However, since our approach is based on the depth map, it primarily reconstructs the front view of the scene.
To address this, we identify boundary points by checking the inequality: $\arcsin(\vec{n}\cdot\vec{v})<\tau$, where $\vec{n}$ is the triangle’s normal, $\vec{v}$ is the view direction, and $\tau$ (default 3 degrees) is an angular threshold. Vertices satisfying this condition are marked as boundary points, which are further projected with the depth of the nearest non-boundary pixel. A final pass removes triangles connecting different objects, which unfortunately leads to white borders around foreground objects. To resolve this, we introduce boundary duplication (BD) mesh reconstruction, where new vertices share the 2D coordinates of the foreground boundary but use background depth, effectively eliminating borders. See Fig. 3 as an example.
3.2 Progressive Differentiable Rendering
Our method assumes ambient lighting, represented by an environment map, and does not explicitly model light sources within the scene. Ignoring emitters in the scene, the rendering equation becomes the reflected radiance equation:
[…]
where $L_{o}(\bm{x},\vec{\omega}_{o})$ is the outgoing radiance at the world space position $\bm{x}$ in the direction $\vec{\omega}_{o}$, $f_{r}(\bm{x},\vec{\omega}_{i},\vec{\omega}_{o})$ is the bidirectional reflectance distribution function (BRDF) at $\bm{x}$, representing the ratio of reflected radiance in the direction $\vec{\omega}_{o}$ to irradiance from the direction $\vec{\omega}_{i}$, while $L_{i}(\bm{x},\vec{\omega}_{i})$ is the incident radiance at $\bm{x}$ from the direction $\vec{\omega}_{i}$, $\vec{n}$ is the surface normal at $\bm{x}$, and $\Omega$ is the hemisphere around $\vec{n}$ covering all directions of incidence. We let $\mathbf{S}(\bm{x})=\bm{x}_{s}$ denote the transformation from world space to screen space coordinates. Detailed steps are in the supplementary material.
Simplified DisneyBRDF.
To model surface scattering we use the DisneyBRDF [8, 9] for our implementation. During the optimization process, we do not model sheen, clearcoat, and glass, so the BRDF in Eq. 4 can be written
$f_{r}(\bm{x},\vec{\omega}_{i},\vec{\omega}_{o})=f_{s}+(1-\mathbf{M}(\bm{x}_{s}))f_{d}$ (5)
with the specular ($f_{s}$) and diffuse ($f_{d}$) terms defined by
$f_{s}$ $=\frac{F_{s}(\vec{\omega}_{i},\vec{\omega}_{o};\eta)D_{s}(\vec{h};\mathbf{R}(\bm{x}_{s}))G_{s}(\vec{\omega}_{i},\vec{\omega_{o}};\mathbf{R}(\bm{x}_{s}))}{4\,|\vec{n}\cdot\vec{\omega}_{i}||\vec{n}\cdot\vec{\omega}_{o}|}$ (6)
$f_{d}$ $=\frac{\mathbf{A}(\bm{x}_{s})}{\pi}F_{d}(\vec{\omega}_{i})F_{d}(\vec{\omega}_{o})\,,$ (7)
where $F$ is the Fresnel reflectance, which here uses the half vector $\vec{h}=(\vec{\omega}_{i}+\vec{\omega}_{o})/\|\vec{\omega}_{i}+\vec{\omega}_{o}\|$ and models the fraction of light reflected from the surface as a function of material properties, including the relative index of refraction $\eta$. We use the modified Schlick approximation [8]:
$\!\!F_{s}(\vec{\omega}_{i},\vec{\omega}_{o};\eta)$ $=C_{0}(\eta)+(1-C_{0}(\eta))(1-\vec{h}\cdot\vec{\omega}_{o})^{5}\,$ (8)
$C_{0}(\eta)$ $=R_{0}(\eta)(1-\mathbf{M}(\bm{x}_{s}))+\mathbf{M}(\bm{x}_{s})\mathbf{A}(\bm{x}_{s})\,,$ (9)
where $R_{0}(\eta)$ is the Fresnel reflectance of a dielectric at normal incidence. Similarly, for the diffuse term,
$F_{d}(\vec{\omega})$ $=1+(F_{D90}-1)(1-\vec{n}\cdot\vec{\omega})^{5}\,$ (10)
$F_{D90}$ $=0.5+2\,\mathbf{R}(\bm{x}_{s})\,(\vec{h}\cdot\vec{\omega})^{2}\,.$ (11)
Regarding the other terms, $D_{s}$ is the microfacet normal distribution function, while $G_{s}$ is the microfacet shadowing-masking function [51]. We use the GGX distribution [61] for these terms due to its simplicity, with just one roughness parameter $\mathbf{R}(\bm{x}_{s})$, while being able to fit empirical data well.
Environment Map Optimization.
We optimize the environment map $\mathbf{E}$ (envmap) using a position-embedded multi-layer perceptron (MLP), similar to the approach used in NeRF [37]. Although optimizing $\mathbf{E}$ directly without using a neural network is theoretically possible, this often yields suboptimal results. Standard MLPs are generally insensitive to high-frequency information, making pixel-level optimization of the image difficult; CNNs encounter similar limitations. The envmap is obtained using
$\mathbf{E}=\mathrm{MLP}_{\text{env}}(\gamma(\mathbf{e});\theta_{\text{env}})\,,$ (12)
where $\gamma(\cdot)$ is a positional encoding function and $\theta_{\text{env}}$ is the set of parameters for the MLP. We use Softplus after the final layer in $\mathrm{MLP}_{\text{env}}$ to ensure that the output has only positive values. The envmap $\mathbf{E}$ and the input image $\mathbf{e}$ are both in $\mathbb{R}^{(w\times h)\times c}$ with the input image being initially noise.
In a differentiable rendering of the geometry $\mathbf{D}_{p},\mathbf{N}_{p}$ based on Monte Carlo integration of Eq. 4, we evaluate the BRDF using $\mathbf{A}_{p},\mathbf{R}_{p},\mathbf{M}_{p}$ (Eq. 5) and the incident radiance using
$L_{i}(\bm{x},\vec{\omega}_{i})=V(\bm{x},\vec{\omega}_{i})L_{\mathbf{E}}(\vec{\omega}_{i})\,,$ (13)
where $L_{\mathbf{E}}(\vec{\omega}_{i})$ is the radiance from the envmap in the direction $\vec{\omega}_{i}$ and $V(\bm{x},\vec{\omega}_{i})$ is visibility, which is zero if geometry is found in the direction $\vec{\omega}_{i}$ and one otherwise.
Our optimization goal is to find the $\mathbf{E}^{\ast}$ that minimizes the difference between the rendering and the ground truth ($I$):
$\mathbf{E}^{\ast}=\operatorname*{\arg\!\min}_{\mathbf{E}}\mathcal{L_{\mathbf{E}}}(L_{o}(\bm{x},\vec{\omega}_{o}),I)$ (14)
where environment map loss $\mathcal{L_{\mathbf{E}}}$ combines $L_{1}$ and $L_{2}$ loss. The network weights $\theta_{\text{env}}$ are discarded after optimization as we only need $\mathbf{E}^{\ast}$.
Material Properties Optimization.
Once we have an optimized envmap we furthur refine the obtained material properties given in section 4.1 using a position-embedded MLP such that :
$\left[\mathbf{A},\mathbf{R},\mathbf{M}\right]=\mathrm{MLP}_{\text{mat}}(\gamma(\mathbf{m});\theta_{\text{mat}})$ (15)
$\mathbf{m}=\left[\mathbf{A}_{p},\mathbf{R}_{p},\mathbf{M}_{p}\right]\,,$ (16)
[…]
We use the following formula as the activation function after the final layer:
$\left[\mathbf{A},\mathbf{R},\mathbf{M}\right]=\zeta\,\tanh(\mathbf{z})+\mathbf{m}\,,$ (17)
where $\mathbf{z}$ is the output from the final layer, and $\zeta$ is a scaling factor. Unlike regular images, $\mathbf{A},\mathbf{R},\mathbf{M}$ may contain many values equal to 0 or 1. Approximating these extremes requires $\tanh$ to output very large or small $\mathbf{z}$ values, which can hinder optimization. The scale factor $\zeta$ mitigates this by adjusting the range of $\mathbf{z}$, facilitating smoother training.
The rendering process is the same as for the envmap, except that now we use $\bm{E}^{\ast}$ and keep it fixed.
Our optimization goal is to find $\mathbf{A}^{\ast},\mathbf{R}^{\ast},\mathbf{M}^{\ast}$ that minimize the difference between the rendered result and the ground truth ($I$):
$\mathbf{A}^{\ast},\mathbf{R}^{\ast},\mathbf{M}^{\ast}=\operatorname*{\arg\!\min}_{\mathbf{A,R,M}}\mathcal{L}_{\text{Mat}}(L_{o}(\bm{x},\vec{\omega}_{o}),I)$ (18)
where $\mathcal{L}_{\text{Mat}}$ is the material loss:
$\mathcal{L}_{\text{Mat}}$ $=\mathcal{L}_{\text{re}}+\delta\mathcal{L}_{\text{cons}}\,,$ (19)
and the individual terms are:
$\mathcal{L}_{\text{re}}$ $=\mathcal{L}_{L1}(L_{o},I)+\mathcal{L}_{L2}(L_{o},I)$ (20)
$\mathcal{L}_{\text{cons}}$ $=\sum_{X\in\{\mathbf{A},\mathbf{R},\mathbf{M}\}}\mathcal{L}_{L1}(X_{o},X_{p})\,,$ (21)
where $\mathcal{L}_{L1}$ and $\mathcal{L}_{L2}$ represent the $L_{1}$ and $L_{2}$ losses, respectively, and $\mathcal{L}_{\text{re}}$ measures the rendering loss between $L_{o}$ and $I$. The constraint loss $\mathcal{L}_{\text{cons}}$ compares the optimized and predicted material properties, introducing constraints to align the optimization with MatNet’s predictions while minimizing the difference between $L_{o}$ and $I$. The scaling factor $\delta$ controls the deviation from MatNet predictions: a smaller $\delta$ allows more flexibility, while a larger $\delta$ keeps the optimization closer to MatNet, potentially increasing the gap between the rendering and the target.
Optimization can be in stages: first optimizing $\mathbf{R}$ and $\mathbf{M}$, then $\mathbf{A}$. As albedo impacts rendering most significantly, increasing its scale factor in $\mathcal{L}_{\text{cons}}$ during joint optimization prevents the network from overfocusing on albedo, ensuring balanced material property optimization. Besides, our method can optimize normals, but we found optimizing only $\mathbf{A}$, $\mathbf{R}$, and $\mathbf{M}$ yields sufficiently good results, while adding normals degrades performance. See supplementary material for details.
3.3 Opaque Material Editing
For $\mathbf{R,M,A}$ edits (excluding transparency), we use the SAM2 [47] model to segment the input image, creating a mask for the edit region. We can then modify values within this region and re-render an image. For $\mathbf{A}$ edits, we convert to HSV color space for intuitive color adjustments.
Material Transparency Editing.
As described in Sec. 3.1, we reconstruct the mesh using the depth map, capturing only the visible front side of objects (see Fig. 3). This limits transparent object editing, as refractive effects cannot be fully modeled. While existing methods generate multi-view images for mesh reconstruction via NeRF [37, 31, 32, 54] or Gaussian Splatting [17, 23, 74], we aim to solve Material Transparency Editing directly within a single-view constraints, avoiding full scene reconstruction. We model transparent objects using DisneyBSDF [9] as:
$f_{\text{glass}}=\frac{\sqrt{\mathbf{A}_{\text{glass}}(\bm{x}_{s})}(1-F_{s})D_{s}G_{s}|\vec{h}\cdot\vec{\omega}_{o}||\vec{h}\cdot\vec{\omega}_{i}|}{|\vec{n}\cdot\vec{\omega}_{i}||\vec{n}\cdot\vec{\omega}_{o}|(\vec{h}\cdot\vec{\omega}_{i}+\eta\,\vec{h}\cdot\vec{\omega}_{o})^{2}}\,,$ (22)
where $\mathbf{A}_{\text{glass}}(\bm{x}_{s})=(1-\mathbf{M}(\bm{x}_{s}))\mathbf{A}_{\text{BG}}(\bm{x}_{s})T$, with $T$ denoting specular transmission and $\mathbf{A}_{\text{BG}}$ the albedo of the diffuse background object after refraction (distinct from $\mathbf{A}$). Thus, by knowing the refraction points, we can simulate transparency effects without a back-side mesh, supporting single-view transparency editing.
We use SAM2 [47] for object mask extraction and an inpainting model [48] to generate $\mathbf{A}_{\text{BG}}$. For an incoming ray of direction $-\vec{\omega}_{o}$ hitting the masked area, the direction of the transmitted ray $\vec{\omega}_{t}$ is computed using the law of refraction [22].
Assuming the object’s back normal is same as its front normal $\mathbf{N}_{p}(\bm{x}_{s})$, the world position after two refractions is
$\bm{x}_{2}=\bm{x}+d_{1}\,\vec{\omega}_{t1}+d_{2}\,\vec{\omega}_{t2}\,,$ (23)
where $\bm{x}$ is the initial position of incidence, $d_{1}$ and $d_{2}$ are refraction distances, and $\vec{\omega}_{t1}$ and $\vec{\omega}_{t2}$ are the directions after first and second refractions, respectively. Note that $\vec{\omega}_{t2}$ depends on $\mathbf{N}_{p}(\mathbf{S}(\bm{x}+d_{1}\,\vec{\omega}_{t1}))$. We trained a neural network on synthetic data to predict $d_{1}$ and $d_{2}$ for various shapes. Thus, $\mathbf{A}_{\text{BG}}(\bm{x}_{s})=\mathbf{A}_{\text{BG}}(\mathbf{S}(\bm{x}_{2}))$, where $\mathbf{S}(\bm{x_{2}})$ are the screen coordinates of $\bm{x}_{2}$ after double refraction.
Refraction Length Prediction.
We briefly explain the synthetic dataset preparation and training strategy, for more details we refer to the supplementary material.
Dataset Preparation: Due to multiple refractions within the object and varying refraction distances depending on the IOR, we simplify the refraction distance for single transmission when the IOR is approximately 1. For each object, we randomly selected 200 camera positions, pairing each with 200 different envmaps as backgrounds.
Training: We normalize the refraction length to [0,1] and use the object along with its corresponding mask as input, with the refraction length as ground truth, following the same training method used for the depth map.
4 Experiments
This section evaluates the performance of our method.
4.1 Material Prediction Accuracy
Synthetic Dataset.
We evaluate the performance of our MatNet on the testset of the InteriorVerse [78] synthetic indoor dataset. As shown in Table 1, our method achieves better prediction accuracy across all evaluated metrics. Qualitative comparisons are shown in Fig. 4.
Table 1: Quantitative Evaluation of Material Prediction on Synthetic Dataset [78].
Albedo Roughness Metallic Depth Normal
[…]
Li et al. [29] 6.173 0.488 0.529 8.665 0.529 0.559 - - - - - 37.696
Zhu et al.[78] 13.407 0.680 0.302 13.670 0.596 0.360 15.600 0.489 0.281 1.828 2.045 26.872
Kocsis et al. [24] 12.774 0.656 0.307 9.070 0.516 0.384 7.228 0.205 0.375 - - -
Ours 19.530 0.811 0.108 16.632 0.663 0.294 18.660 0.701 0.243 0.244 0.491 11.359
Figure 4: Qualitative results of material prediction on synthetic dataset.
Real World Dataset.
Figure 5: Qualitative results of material prediction on the real world dataset IIW [4]. Zhu et al. [78] retained too many details from the original image while Kocsis et al. [24], due to their reliance on Stable Diffusion, altered image details excessively. Additionally, Kocsis et al.’s method struggles to handle specular highlights and facial features effectively.
Table 2: Quantitative evaluation, WHDR ($\delta=0.1$), of albedo predictions on the IIWild Dataset [4].
WHDR $\downarrow$
Li et al. [29] 0.342
Zhu et al. [78] 0.232
Kocsis et al. [24] 0.206
Ours 0.197
We assess the performance of albedo predictions using the IIW dataset [4], which includes sparse pairwise human annotations for albedo. For evaluation, we use the metric developed for the IIW dataset, Weighted Human Disagreement Rate (WHDR), which quantifies the error when predicted albedo values differ from human-provided annotations. As shown in Table 2, our model achieves better results with a lower WHDR score. Fig. 5 provides a qualitative comparison.
4.2 Environment Map Optimization
Currently, single-view envmap generation primarily relies on Stable Diffusion [44], which can only produce visually plausible envmaps. Optimization-based methods often require multi-view input [60, 72, 42] or use implicit lighting representations [69, 78, 24]. For optimization approaches, we selected DPI [35] as our benchmark, adapting it to single-view input and providing it with a mesh reconstructed using our method (Sec. 3.1).
Table 3: Quantitative evaluation of environment map optimization. Envmap measures the difference between the optimized and ground truth envmaps, while Rerender represents the difference between the rendered relit scene with the optimized envmap and the ground truth. SH ERROR calculates the error as the norm of the difference matrix between the spherical harmonics of the ground truth and optimized envmaps. The label mat indicates the material properties predicted by MatNet.
[…]
Envmap Rerender Envmap Rerender Envmap Rerender Envmap
DiffusionLight [44] 0.279 0.342 0.635 13.595 8.728 0.134 7.800
DPI [35] 0.403 0.640 1.115 15.155 7.767 0.166 6.768
Ours w/o mat 0.414 0.518 0.902 10.242 8.302 0.355 5.571
Ours w/ mat 0.703 0.784 4.507 18.981 6.734 0.023 3.049
As our focus is not high-resolution envmaps, we limit the resolution to $32\times 16$, consistent with prior work [28, 12]. We evaluated three scenes, each with 60 envmap optimization results. Evaluation metrics included the difference between the optimized and the ground truth envmaps and the difference between relit scenes using the optimized and the ground truth envmaps. Both methods used ground truth material properties for relighting. As shown in Table 1, our method outperformed the others across all metrics, benefiting from the MatNet-predicted properties $\mathbf{A}_{p},\mathbf{R}_{p},\mathbf{M}_{p}$ during envmap optimization. Qualitative results are in Fig. 6.
Figure 6: Qualitative results of envmap optimization.
Notably, while rerendering error is low, envmap error remains relatively high due to the inherent ambiguity in single-view optimization; accurate lighting in certain regions can yield a close match to the ground truth, even if the overall envmap differs significantly.
4.3 Object insertion and relighting
Since our rendering process is physically based, it enables the rendering of realistic shadows in object insertion tasks, which significantly enhances the realism of image editing, as seen in Fig. 7.
Figure 7: Qualitative results of object insertion and relighting, compared with the multi-view method MAIR (Choi et al. [12]) and the single-view method by Li et al. [27]. Note that relight result2 has a realistic shadow.
Additionally, to validate the visual accuracy of our approach, we used real photograph comparisons instead of synthetic data. We 3D scanned a glass jug to obtain its mesh and took photographs of the scene from the same angle with and without the jug. We then inserted the mesh jug into the image without the jug and compared it with the actual photograph, as demonstrated in Fig. 8.
Since single-view inverse rendering mostly uses neural networks for rendering and does not support object insertion of transparent objects, we only compare the difference between ground truth and our method.
Figure 8: Comparison between ground truth (GT) and our object insertion result for real world image. The error in the reflective area mainly comes from the insufficient accuracy of optimized envmap, as explained in Sec. 4.2.
4.4 Material Editing
For transparency editing evaluation, we used real-world photo comparisons. A 3D-printed opaque jug, identical in size to the scanned mesh from Sec. 4.3, was placed alongside a glass jug at a fixed camera position, and two photos were captured as ground truth for transparency editing. Since Alchemist [53], closely related to our work, is not open-sourced, we compared our method with iPix2Pix [7] and Stable Diffusion inpainting [48]. For iPix2Pix, text prompts were used for editing, while Stable Diffusion inpainting received both a mask of the jug region and a text prompt. Results are shown in Fig. 9.
Figure 9: Material transparency editing compared with iPix2Pix [7] and Stable Diffusion inpainting [48]. Our method can more accurately replicate the distortion effects caused by light refraction.
Figure 10: Qualitative evaluation of material editing. Our method enables albedo editing to various colors, whereas Alchemist [53] is limited to white. For roughness editing, our approach follows physical principles—when roughness is set to 0.2, a realistic reflection of the stone appears at the bottom of the bottle. Our method preserves Charizard’s color during metallic edits and shows that increasing metallic alone darkens the material. In transparency editing, our approach accurately captures light refraction and maintains the vase’s appearance when transparency is increased.
For albedo, roughness, and metallic evaluation, since Alchemist [53] is not open-sourced, we extracted images from its paper and applied our material editing for comparison. These images are notably out-of-domain for indoor scenes. While MatNet predicts accurately for indoor scenes, only albedo predictions remain reliable, with roughness and metallic predictions being less consistent. However, differential rendering allows us to perform inverse rendering on real photos using only albedo. To compensate for unreliable roughness and metallic predictions, we manually set priors for each image to improve optimization. Results are shown in Fig. 10. Detailed settings and user study are provided in the supplementary material.
4.5 Human Face Relighting
We also tested the performance of our method on human face relighting to demonstrate its potential. We used Material Net, trained on the indoor dataset and not specifically optimized for the human face dataset, to predict the $\mathbf{A},\mathbf{R},\mathbf{M}$, followed by progressive optimization. The optimized results were used to relit. It is important to note that, despite the seemingly good results (as shown in Figure 11), we did not model the subsurface scattering, which is critical for realistic face rendering. In our approach, subsurface scattering component is fixed within the albedo during optimization and does not change with lighting, thus making the rendered results inaccurate. This could be possibly improved in future work by integrating subsurface scattering into the rendering pipeline. Additionally, if there are specular highlights on the face, they too would be fixed in the albedo during the optimization process.
Figure 11: Our method was applied to relight real human faces (FFHQ). When the face lacks strong highlights, the relighting results are more accurate (above). However, in cases where highlights are present (below), these highlights tend to be baked into the albedo during optimization, causing them to persist even after the lighting conditions are altered.
5 Ablations
We tested the results of envmap optimization under two conditions: providing MatNet predictions versus setting $\mathbf{A}=\mathbf{R}=0.5,\mathbf{M}=0.1$. As shown in Table 3, providing MatNet predictions significantly improves the accuracy of envmap optimization.
We also conducted an ablation study on the value of $\delta$ in Eq. 19. Using 20 randomly selected images from the IIW dataset [4], we tested different $\delta$ values and compared the rerendered images with the input images. We evaluated two strategies: simultaneous optimization of $\mathbf{ARM}$ and sequential optimization ($\mathbf{RM\&A}$), where $\mathbf{RM}$ is optimized first, followed by $\mathbf{A}$. To save time, we applied early stopping, halting optimization if $\mathcal{L}_{\text{re}}$ reduced by less than 5% over 20 consecutive steps. Results are shown in Table 4. For additional studies, see the supplementary material.
Table 4: Ablation of $\delta$ in Eq. 19. Comparing the rerender error.
Opt $\mathbf{ARM}$ Opt $\mathbf{RM}$&$\mathbf{A}$
$\delta$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$
0.0 0.459 13.479 0.0504 0.611 17.390 0.0342
0.5 0.473 14.490 0.0473 0.699 18.242 0.0271
1.0 0.542 15.363 0.0426 0.618 17.396 0.0312
1.5 0.554 15.403 0.0446 0.637 17.541 0.0313
2.0 0.571 15.320 0.0424 0.640 17.719 0.0280
2.5 0.610 16.151 0.0371 0.640 16.922 0.0304
3.0 0.612 16.109 0.0366 0.656 16.819 0.0299
6 Limitations and Discussion
Figure 12: Super-resolution can improve image quality while preserving overall structure, such as lighting and shadows. However, Stable Diffusion-based super-resolution may alter some details, as seen in the excessive modifications to the statue in the right image.
Since our images are rendered and mesh reconstruction limitations can cause artifacts along object edges, using super-resolution (SR) [48, 59, 65, 66] may help reduce these artifacts. However, SR models often overly modify the image, as shown in Fig. 12. Improving SR models by giving additional input could potentially mitigate this issue.
Finally, physically based in this paper refers to physically based rendering (PBR), a computer graphics approach that may not strictly follow physical laws. For more general limitations, please refer to the supplementary material.
7 Conclusion
In this work, we present a novel pipeline for physically-based inverse rendering that combines neural network predictions with differentiable rendering to optimize material properties and lighting conditions from a single image. Our approach enables accurate relighting, transparent object insertion, and material editing tasks, including transparency editing without requiring complete geometry.
Supplementary Material
8 More Ablations
Here we provide more ablation results.
8.1 Optimize SH coefficients instead of envmap
We tested using SH coefficients instead of envmaps and found that optimizing SH coefficients took longer to reach the same loss. We use 3rd-order SH coefficients to represent the light source. Given the ground truth material properties, we perform light source optimization under different lighting conditions. The time taken by each method to achieve the same $\mathcal{L}_{\text{re}}$ is compared. The results is shown if Table 5.
Table 5: Time taken by different methods to optimize $\mathcal{L}_{\text{re}}$ to 0.03. Direct envmap optimization is faster than optimize SH coefficient.
SH coeff Envmap
Time 473 s 103 s
8.2 Optimization Steps and Rerendering error
We analyzed the relationship between optimization steps and rerender error. We randomly selected 10 images from the InteriorVerse dataset [78] to test rerendering error at different optimization steps. Optimization on an RTX 3090 runs at 2 steps per second. Based on MatNet’s accuracy, satisfactory results are typically achieved after 10 minutes of optimization. Results are shown in Table. 6.
Table 6: Ablation of Optimization steps. Compare the rerender error.
Opt Steps SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$
50 0.645 15.529 0.0180
100 0.750 18.735 0.0134
1000 0.825 25.675 0.0034
2000 0.876 28.204 0.0015
8.3 Optimization Network
We experimented with using a CNN-based UNet for Material Properties optimization. The UNet architecture includes 2 downsampling blocks and 2 upsampling blocks, with each block containing two 2D convolution layers (kernel size = 3, padding = 1). Additionally, we compared direct optimization of material properties without a neural network. As shown in Fig. 13, direct optimization was the slowest, followed by the CNN-based UNet, while the position-embedded MLP achieved the fastest optimization speed.
Figure 13: Compared to a CNN-based UNet, the position-embedded MLP achieves faster optimization.
Additionally, due to the characteristics of convolutional neural networks, noticeable artifacts may appear during optimization. As shown in Fig. 14, the optimized images show a regular pattern of black dots.
Figure 14: UNet Artifect. The red box highlights an enlarged view of this area, and the details is in the second row.
It is worth noting that when MatNet predictions are unreliable (e.g., for roughness and metallic), directly optimizing material properties without relying on the neural network can achieve lower $\mathcal{L}_{\text{re}}$ given sufficient optimization time (such as more than 3000 optimization steps). Table 7 shows the minimum MSE render loss achievable by both methods in the absence of roughness and metallic predictions.
Table 7: Comparison of the lowest $\mathcal{L}_{\text{re}}$ that can be obtained using MLP and directly optimizing material properties when the predicted roughness and metallic are not provided.
w/o Network w/ MLP
Min $\mathcal{L}_{\text{re}}$ 0.0038 0.0087
8.4 Optimizing Normal
As shown in Fig. 15, optimizing material properties while simultaneously optimizing normals often leads to convergence difficulties. To address this, directly optimizing normals without relying on a neural network can help mitigate these issues.
Figure 15: Optimizing normal at the same time will make the optimization process more unstable.
9 Limitations
In this section, we discuss potential limitations of our proposed method from a broader, user-oriented perspective. It is important to note that some points, such as the need for per-image optimization, are not typically viewed as limitations within the differentiable rendering field. However, from a user’s viewpoint, this may make our approach less convenient than stable diffusion-based methods [53, 7].
1. As with all differentiable rendering methods, our approach requires optimization for each image to obtain the lighting and optimized material properties, which might be quite time-consuming. Depending on MatNet’s prediction accuracy, optimization takes between 5 to 30 minutes. For out-of-domain images (i.e., non-indoor scenes) or higher-quality requirements, optimization time may increase. This limits our method’s ability to achieve the rapid, batch image editing enabled by stable diffusion-based techniques [7]. Training MatNet on a broader dataset for more accurate predictions could help mitigate this limitation.
2. Our method uses an envmap to represent lighting, complicating accurate modeling of physical light sources in the scene. If a light source is present in the original image, optimization may fix it within the albedo, leading to inaccuracies during relighting.
3. The introduction of differentiable rendering allows our method to perform well even on out-of-domain images. However, if MatNet predictions are poor, differential rendering optimization requires careful tuning for each image, such as adding a mask to the object to be edited or experimenting with different optimization strategies.
4. For Material Transparency Editing, complex object geometry can lead to inaccurate refraction distortions, as our method assumes only two refractions. Single-view mesh reconstruction [31, 32, 54] may yield more accurate refractions for complex shapes. Besides, since we simulate refraction without the rays truly passing through the object, shadows cast by transparent objects may lack accuracy, and does not have caustic effect. (Note that this problem does not exist for object insertion tasks with complete geometry.)
5. Due to mesh reconstruction limitations, artifacts may appear along object edges during strong relighting, as the mesh is discontinuous at these edges. This can be mitigated using super resolution.
6. For images with transparent objects, super-resolution often yields unreliable results and may require multiple samples to reach an acceptable outcome, as shown in Fig. 16. (This is not a limitation of our method, but is noted here due to the potential need for SR techniques.)
Figure 16: Super-resolution often struggles with refraction effects. As shown, both Super Resolution 1 and Super Resolution 2 exhibit noticeable refraction issues, requiring multiple runs to achieve a satisfactory result, as seen in Super Resolution 3.
10 Refraction Length Prediction
Refraction plays a critical role in visual accuracy in physically based rendering. Our pipeline is fine tuned for single view material and geometric estimation. This however means that information behind a reflective/refractive interface is intractable from a single view without any additional geometric information. Furthermore, this also means that objects cannot be edited accurately when secondary interfaces are crucial to the visual fidelity. To overcome this issue we train another neural network essential to predict secondary refractive interfaces if a user wants to change any material to transparent object such as glass, water etc. We describe this pipeline in this section.
10.1 Dataset Preparation
To create a dataset to predict the secondary interface distance, we collected a set of $200$ environment maps, $50$ triangle meshes of objects commonly found in the indoors environments like cups, fruits etc. Next we predefined a set of $200$ viewpoints that generally covers the entire exterior surface of an object.
The data collection was done in 2 parts, first the set of $200$ environment maps were rendered without any object in the scene which illustrates the background of the scene. This dataset was $200$ (Environment maps) $\times$ $200$ (Viewpoints) images in quantity. The second part of the data collection was done in scenes without an environment map, where the objects were placed in the center of the scene and the $200$ viewpoints were used to capture the distance to the secondary interface if the object were to be transparent. The distances can be generally defined as,
$d=\begin{cases}\|\mathcal{P}_{n+1}-\mathcal{P}_{n}\|,&\text{if }\vec{\omega}_{r}\cdot\vec{n}>0\text{ \& hollow object}\\ \|\mathcal{P}_{n+1}-\mathcal{P}_{n}\|,&\text{if }\vec{\omega}_{r}\cdot\vec{n}<0\text{ \& non-hollow object}\\ 0,&\text{Otherwise }\end{cases}$ (24)
where,
  • $\mathcal{P}_{n}$ : is the $n$-th intersection point in 3D world coordinates,
  • $\vec{\omega}_{r}\cdot\vec{n}$ : is the cosine term of the ray direction $\vec{\omega}_{r}$ and the normal $\vec{n}$.
Furthermore, "hollow" objects represent things like water jugs with air in between the surfaces and "non-hollow" objects are things like lemon with no air inside the object.
Examples are shown in Fig. 17.
Figure 17: Example of refraction length dataset. The closer to white, the longer the refraction distance.
11 Material Editing Settings
The selected images are out-of-domain for the indoor dataset, resulting in MatNet predictions where only albedo is relatively accurate, while roughness and metallic predictions are poor and unsuitable as initial conditions for optimization. Therefore, specific optimization conditions must be set for each image to achieve satisfactory results.
Cherry. The predicted roughness and metallic are relatively inaccurate, leading to excessive red light in the envmap if directly used for optimization. To address this, we initialize the envmap as pure white and start with material property optimization. The default initial values for roughness and metallic are set to 0.5. After optimizing roughness and metallic, we proceed to envmap optimization, followed by albedo optimization, as albedo predictions are more reliable.
Bottle. MatNet predictions are relatively accurate, and optimization follows the method described in the main paper.
Charizard and Vase. Similar to Cherry, we initialize the envmap as pure white and start with material property optimization. Roughness and metallic are initialized to 0.5, while albedo uses MatNet predictions. Optimization begins with roughness and metallic, followed by envmap optimization, and finally albedo optimization. During editing, setting all mask values uniformly for metallic without adjusting roughness can introduce artifacts. To improve material editing results, we use a mask during optimization. Specifically, SAM2 is used to segment Charizard’s and vase’s mask, roughness and metallic values are unified within the mask during optimization. This ensures better editing results.
12 Material Editing User Study
We conducted a user study on material editing results, asking participants to rank the editing results from Alchemist [53], ours (direct rendering), and ours SR (rendered result refined with Super Resolution). The highest rank received 3 points, and the lowest 1 point. Since these are real-world photos with no ground truth, participants were provided with synthetic scene renderings as reference and asked to imagine the expected edit on real photos as closely as possible. It should be noted that participants generally found it challenging to envision the correct editing result, making it difficult to judge which method was better. When uncertain, participants often zoomed in to examine details for ranking, which may have led to lower scores for our direct rendering. However, our scores improved significantly after applying super resolution.
Table 8: Material Editing User Study. SR represent for super resolution.
Edit Task Alchemist [53] Ours Ours (SR)
Albedo 1.04 2.36 2.61
Roughness 2.21 1.57 2.21
Metallic 2.50 1.79 1.71
Trans 1.93 1.64 2.43
Sum 7.68 7.36 8.96
Figure 18: Our methods with super-resolution. Using super resolution (SR) can greatly improve the quality of the rendered image. However, the current SR model cannot fully retain the details of the original image and will make excessive modifications to the original image. We also provide the effect of modifications to material properties on appearance in a synthetic dataset.
13 Coordinate Transformation
Since our input into the renderer are based in screen space, we briefly describe how to transform a world-position $\bm{x}$ into screen-space. Extending $\bm{x}$ into homogeneous coordinates $\bm{W}=[x,y,z,1]^{T}$, we first transform into camera space coordinates by
$\mathbf{Q}=E^{-1}\cdot\bm{W}$ (25)
where $E$ is the extrinsic camera matrix. For a perspective projection matrix $\mathbf{P}$ we can obtain the Normalized Device Coordinates (NDC) by:
$\mathbf{N}=\bm{P}\cdot\mathbf{Q}$ (26)
Converting $\mathbf{N}$ into inhomogeneous coordinates we can obtain the screen-space coordinates:
$x_{\text{screen}}=\left(\frac{x_{\text{ndc}}+1}{2}\right)\cdot W$ (27)
$y_{\text{screen}}=\left(\frac{y_{\text{ndc}}+1}{2}\right)\cdot H$
This process maps the world position to a specific location on the screen which we will denotes as $\bm{x}_{s}=\mathbf{S}(\bm{x})=[x_{\text{screen}},y_{\text{screen}}]$. In the following $\mathbf{A}$, $\mathbf{R}$ and $\mathbf{M}$ represents the image texture values of albedo, roughness and metallic. Thus for a shading point $\bm{x}$, we can obtain the material properties with screen coordinate indexing i.e. $\mathbf{A}(\bm{x}_{s}),\mathbf{R}(\bm{x}_{s})$ and $\mathbf{M}(\bm{x}_{s})$.